OpenStudy (alante):

Using synthetic division, find (x4 − 16x2) (x − 4) whats the answer???

OpenStudy (anonymous):

(x4−16x2)(x−4) =(x4+−16x2)(x+−4) =(x4)(x)+(x4)(−4)+(−16x2)(x)+(−16x2)(−4) =x5−4x4−16x3+64x2